Questions 6-12 Zestas Economic Data 2016 | 2017 Actual GDP ($ Billions) 900 B = ? = Actual GDP growth rate (%) 2.0% Real GDPGDP in cell A (rounded to closest 10 dollars) is a) 854 b) 866 c) 890 d) 905 GDP in cell B (rounded to the closest 10 dollarsHello, can someone help me out with the question 8,9,11,12?

I calculated for question 8 the answer might be 0.33, however, there is no 0.33 as the answer.

The procedure is (Real GDP2017-Real GDP2016)/ Real GDP2016=(893-890)/890=0.33

Thank you so much!
Answer #1

8. The correct option would be

  • (a) 0.01

Since no rounding is mentioned, using the original non-rounded values, we have the following values. 900 RGDP2016 = = 891.0891 , NGD P2017 = 900*(1 +0.02) = 918 , _NGD 2017 _ 918 KGDF 2017 = 1.031.03 $41.2021 , and hence RGD P2017 - RGDP 2016 G 2017 = ? RGDP 2016 or G 2017 = 891.2621 - 3 * 10C 891.0891 or GR2017 = 0.019414445 = 0.01% (note that this value is not rounded, and is taken upto two decimals).

9. The correct option would be

  • (b) 2.66

We have _RGD P2018 - RGD 2017 G 2018 = + RGDP 2017 or 915 – 891.2021, 100- Gr2018 = 2 891.2621 or GR2018 = 2.663402831 = 2.66% (note that this value is also not rounded, and is taken upto two decimals).

11. The correct option would be

  • (b) 2.10

We have the required values as follows. We have NGDP2018 = NGDP2017 *(1 +0.025) =918(1+0.025) = 940.95 , DEFL2018 = RGDP 2018 NGDP2018 **100= 940.95 915 * 100 = 102.830 , and hence the inflation would be as DEFL2019 - DEFL2018 100 12019 = ? DEFL 2015 105 - 102.836 102.836 - 100 or 12019= 2.104321444 = 2.10% (taken upto two decimals).

12. The correct option would be

  • (d) The AD curve . . . not shift while SRAS and LRAS . . . increase.

A much larger Canadian and world wheat crop than forecast would mean a positive supply shock, which would increase the SRAS and LRAS, since now more can be supplied than before at the same price level.